OpenStudy (anonymous):

The amount of money that is left in a medical savings account is expressed by the equation y=-24x+379, where x represents the number of weeks and y represents the amount of money, in dollars, that is left in the account. After how many weeks will the account have $67 left in it? a)10 weeks b)13 weeks c)15 weeks d)21 weeks

OpenStudy (bee_see):

replace x with the choices they give you.

OpenStudy (bee_see):

A) x=10, -24(x)+379= -24(10)+379=???

OpenStudy (bee_see):

B) x=13 C) x=15 and D)x=21. Which one gives you $67?

OpenStudy (anonymous):

i got i now thanks